PLEASE HELP WITH THIS QUESTION.
Joshua currently earns $5000 a year from his lawn care business. He has two plans to increase his revenue.
Plan 1 would increase his yearly revenue by 20%.
Plan 2 would increase his revenue by $1500 a year.
After 7 years, how much more revenue does Plan 1 generate than Plan 2? Round to the nearest dollar.

Answer by rothauserc(4718)

revenue or interest = Prt where P is the principal, r is rate and t is time
we know plan 1 yields 20% revenue and
plan 2 yields $1500 revenue which is 1500/5000 = 30%, therefore
plan 1 = 5000*.20*7 = $7000
plan 2 = 5000*.30*7 = $10,500
plan 2 generates $3,500 more revenue than plan 1